# RoCoFT: Efficient Finetuning of Large Language Models with Row-Column Updates

### v2

@@ 4 Experiments
Table 2: Results of DeBERTaV3-bas on SQuAD v1.1, v2.0 benchmarks, reported using EM/F1 scores and BART-large on XSum and CNN/Daily Mail, reported using ROUGE metrics (ROUGE-1, ROUGE-2, ROUGE-L).
RoCoFTRow 13.37M 66.33 74.53 73.56 56.60 72.14 73.29 57.48 72.92 79.76 70.94 70.95 70.90 54.42
RoCoFTColumn 13.37M 66.34 74.64 73.12 55.93 72.50 73.11 57.19 72.90 79.72 71.05 70.88 70.76 54.38
RoCoFTRow 11.62M 65.92 68.53 69.90 45.97 66.87 64.91 45.12 65.07 89.45 72.80 80.45 82.12 56.79
RoCoFTColumn 11.62M 65.12 68.22 69.96 45.98 66.78 64.89 45.70 64.81 89.74 72.24 80.23 82.61 56.70
RoCoFTRow 13.47M 69.36 80.01 78.09 89.28 76.73 76.46 60.55 76.96 90.55 77.37 86.12 82.66 60.75
RoCoFTColumn 13.47M 69.32 80.08 77.99 89.46 76.41 76.46 60.59 76.90 90.42 77.35 86.16 82.48 60.35
RoCoFTRow 24.88M 71.46 83.32 79.54 91.86 83.22 83.65 67.12 81.54 90.69 79.70 88.24 87.28 66.60
RoCoFTColumn 24.88M 71.44 83.52 79.50 91.84 83.20 83.39 67.06 81.73 91.46 79.63 88.11 87.58 66.63
Performance Analysis: Table 1 presents the validation performance of RoCoFT compared with baselines on the GLUE benchmark tasks (Wang et al., 2018). RoCoFT achieves competitive or superior results while updating significantly fewer parameters. For instance, RoCoFT3-Row, with only 0.418 million trainable parameters, attains an average score of 85.65/90.61 on RoBERTa-base (Liu et al., 2019), outperforming methods like LoRA (Hu et al., 2021) and MAM Adapter (He et al., 2021), which utilize more parameters. Moreover, RoCoFT variants consistently rank among the top performers across multiple tasks such as the MRPC, QNLI, and RTE, demonstrating robustness and versatility. On RoBERTa-large, RoCoFT3-Row matches the highest MCC score of 67.39 on the CoLA, achieves an accuracy of 96.69% on the SST-2, and attains better performance on MRPC, Quora Question Pairs (QQP), QNLI, and RTE.
As shown in Table 2, our proposed methods demonstrate superior performance on both question answering and summarization tasks while utilizing significantly fewer trainable parameters. Specifically, on the SQuAD v1.1 dataset (Rajpurkar et al., 2016), the RoCoFT3-Row method achieves the highest Exact Match (EM) and F1 scores of 81.70/88.15, outperforming other PEFT methods such as LoRA and AdaLoRA (Zhang et al., 2023b), which require more parameters. Similarly, on SQuAD v2.0 (Rajpurkar et al., 2018), the RoCoFT3-Column attains the top ROUGE-2 score of 8.54 on XSum, showcasing its effectiveness in handling text summarization.
These results underscore RoCoFT’s ability to deliver state-of-the-art performance while maintaining parameter efficiency, making it highly suitable for deployment in resource-constrained environments.
Methods Space Time TTPs APs
FT $O(d\times d)$ $O(d\times d)$ $d^{2}$ 00
$(IA)^{3}$ $O(l_{k}+l_{v}+l_{ff})$ $O(d_{k}+d_{v}+d_{ff})$ $3d$ $3d$
Prompt $O(d\times l_{p})$ $O(d\times l_{p})$ $l_{p}.d$ $l_{p}.d$
LoRA $O((d+d)\times r)$ $O((d+d)\times r)$ $2dr$ $2dr$
LoRA-FA $O((d+d)\times r)$ $O((d+d)\times r)$ $dr$ $2dr$
AdaLoRA $O((d+d+r)\times r)$ $O((d+d+r)\times r)$ $2dr+r^{2}$ $2dr+r^{2}$
LoHA $O(2r\times(d+d))$ $O(2r\times(d+d))$ $4dr$ $4dr$
RoCoFTRow $O(d\times r)$ $O(d\times r)$ $rd$ 00
RoCoFTColumn $O(d\times r)$ $O(d\times r)$ $rd$ 00
Table 4: Space/Time Complexity; Total Trainable Parameters (TTPs) and Additional Parameters in model (Aps) for RoCoFT method and baseline methods for single layer $\mathbf{W}\in\mathbb{R}^{d\times d}$. Within this table, we define $l_{k},l_{v},$ and $l_{ff}$ as the dimensions of three learned vectors in IA3; and $l_{p}$ as the length of the prompt added to the input/layers in prompt tuning and prefix-tuning. For LoRA-type methods, we use $r$ to represent the rank dimension.
Figure 2: Comparison of memory costs for PEFT Methods. Blue bars show the memory cost of the original model weights, while green bars show the memory cost for optimization in each method.
Figure 3: Training time (minutes) comparison across different PEFT methods.
For instance, as shown in Table 1, RoCoFTRow uses only 0.083 million trainable parameters for rank one and 0.418 million for rank three on the GLUE benchmark (Wang et al., 2018), outperforming methods like LoRA (Hu et al., 2021) and MAM Adapter (He et al., 2021), which use 0.89 million and 1.78 million parameters, respectively. Similarly, in question answering and summarization tasks (Table 2), our Row and Column methods utilize just 0.161 million trainable parameters, significantly less than LoRA and AdaLoRA (Zhang et al., 2023b), yet achieve higher or comparable performance.
In terms of computational efficiency (Table 4), our method exhibits lower space and time complexity. Specifically, RoCoFT has a time/space complexity of $O(r\times d)$, compared to LoRA’s $O(2d\times r)$ and Prefix-Tuning’s $O(L\times d\times l_{p})$, where $r$ is the rank, $d$ is the model dimension, $L$ is the number of layers, and $l_{p}$ is the length of the prefix. Moreover, our method does not introduce any additional parameters into the model architecture, which also reduces the total number of parameters and requires less GPU memory and training time, as illustrated in Figure 2. RoCoFT variants have lower memory occupancy during training (approximately $2.85\mathrm{GB}$) compared to other methods like LoRA and AdaLoRA, and consistently require less training time across various datasets, as shown in Figure 3. These results underscore the efficiency of our approach in terms of both parameter count and computational resources, highlighting its suitability for deployment in resource-constrained environments.
5 Finetuning through Neural Tangent Kernel Regression
Kernel methods are classic machine learning algorithms that make use of kernel functions for learning nonlinear mapping of inputs, with SVMs (Cortes & Vapnik, 1995) and Gaussian Processes (Williams & Rasmussen, 2006) being the prime examples. A kernel function $\mathbf{K}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ is a similarity function on the input space $\mathcal{X}$ that satisfies certain symmetry and positive semi-definiteness conditions. If these conditions (Mercer’s conditions) are satisfied the kernel behaves like an inner product over a possibly infinite dimensional space called the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$. The RKHS $\mathcal{H}$ contains functions $f:\mathcal{X}\rightarrow\mathbb{R}$ that maps inputs to real numbers. Although the functions $f\in\mathcal{H}$ are infinite dimensional, when minimizing a loss function $\mathcal{L}$ over a finite training sample $\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{n}$, the Representer Theorem (Schölkopf et al., 2001) tells us that the optimal solution takes the form
$f^{*}(\cdot)=\sum_{i=1}^{n}\alpha_{i}\mathbf{K}(\mathbf{x}_{i},\cdot),$
where $\alpha_{i}$ are coefficients that can depend on labels $\mathbf{y}_{i}$. When the loss function $\mathcal{L}$ is convex the minimization problem is also convex. The ability to learn nonlinear mappings of inputs through solving convex optimization problems makes kernel methods a powerful tool in machine learning.
Kernel methods differ from deep learning with neural networks in that the kernels (and hence the feature representations) are fixed during learning, while deep neural networks continuously update their feature representations during backpropagation. Jacot et al. (2018) made the important discovery that under certain conditions, in the infinite width limit, the training of deep neural networks can be described by a fixed kernel called the Neural Tangent Kernel (NTK). For a neural network function $f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathbb{R}^{k}$ parameterized by $\bm{\theta}$, its Neural Tangent Kernel is defined by
$\mathbf{K}_{{\bm{\theta}}}(\mathbf{x},\mathbf{x}^{\prime})=\langle\nabla f_{{\bm{\theta}}}(\mathbf{x}),\nabla f_{{\bm{\theta}}}(\mathbf{x}^{\prime})\rangle,$
where $\nabla f_{{\bm{\theta}}}(\mathbf{x})$ is the corresponding Jacobian.($\nabla f_{{\bm{\theta}}}(\mathbf{x})$ is $p\times k$ if $\bm{\theta}$ has $p$ parameters, and $\mathbf{K}_{{\bm{\theta}}}(\mathbf{x},\mathbf{x}^{\prime})$ is $k\times k$) This connection allows us to study the behaviour of neural networks with kernel methods. Malladi et al. (2023) extends the NTK theory to model the finetuning of LLMs. As an alternative to finetuning by SGD, given training data $(\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n}$ for a downstream classification task, we can instead solve the following kernel logistic regression problem
where $\mathcal{H}$ is the RKHS defined by the NTK $\mathbf{K}_{\bm{\theta}}$, and $\mathcal{L}(\cdot,\cdot)$ is the logistic loss. For a two-class problem with $\mathbf{y}_{i}\in\{0,1\}$, this is equivalent to
$\min_{\bm{\alpha}}-\sum_{i=1}^{n}\sum_{j=1}^{n}\mathbf{y}_{i}\alpha_{j}\mathbf{K}_{\bm{\theta}}(\mathbf{x}_{i},\mathbf{x}_{j})+\sum_{i=1}^{n}\log(1+\exp(\sum_{j=1}^{n}\alpha_{j}\mathbf{K}_{\bm{\theta}}(\mathbf{x}_{i},\mathbf{x}_{j})))+\frac{\lambda}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{i}\alpha_{j}\mathbf{K}_{\bm{\theta}}(\mathbf{x}_{i},\mathbf{x}_{j}).$
This problem is convex in $\bm{\alpha}$ so the solution is easy to describe without local minima. It is also clear that the solution is completely determined by the value of the NTK $\mathbf{K}_{\bm{\theta}}$ between all training samples $\mathbf{x}_{i}$. Notice that $\bm{\theta}$ is fixed here (usually set to pretrained model weights), so the kernel $\mathbf{K}_{\bm{\theta}}$ is also fixed. Malladi et al. (2023) provides theoretical analysis on conditions when finetuning with SGD will converge to this particular kernel logistic regression setup at the infinite width limit.
Under this framework it becomes feasible to compare full finetuning with finetuning over a subset of parameters by comparing their respective NTKs. Below, we compare the kernels of the 1-row and 1-column version of our RoCoFT method, and we denote the associated parameters as $\bm{\theta}_{{}_{R}},\bm{\theta}_{{}_{C}}\subseteq\bm{\theta}$. The corresponding kernels are defined as
$\mathbf{K}_{{\bm{\theta}_{{}_{R}}}}\!(\mathbf{x},\mathbf{x}^{\prime})=\langle\nabla f_{{\bm{\theta}_{{}_{R}}}}\!(\mathbf{x}),\nabla f_{{\bm{\theta}_{{}_{R}}}}\!(\mathbf{x}^{\prime})\rangle\quad\mbox{and}\quad\mathbf{K}_{{\bm{\theta}_{C}}}\!(\mathbf{x},\mathbf{x}^{\prime})=\langle\nabla f_{{\bm{\theta}_{C}}}\!(\mathbf{x}),\nabla f_{{\bm{\theta}_{C}}}\!(\mathbf{x}^{\prime})\rangle$
Note that while the Jacobians $\nabla f_{\bm{\theta}_{{}_{R}}}\!(\mathbf{x})$ and $\nabla f_{\bm{\theta}_{C}}\!(\mathbf{x})$ can have different dimensions due to different number of parameters, the kernels $\mathbf{K}_{\bm{\theta}_{R}}$ and $\mathbf{K}_{\bm{\theta}_{C}}$ reside in the same function space $\mathcal{H}$ (so does the full finetuning kernel $\mathbf{K}_{\bm{\theta}}$) and can be compared on a data sample.
We first compare few-shot learning performance of these kernels using kernel logistic regression with prompt-based finetuning, as done in Malladi et al. (2023). The kernels are computed with the pretrained RoBERTa-base model. From Table 5 we can see the performance of kernel logistic regression using $\mathbf{K}_{\bm{\theta}_{R}}$ and $\mathbf{K}_{\bm{\theta}_{C}}$ are surprisingly close to using the kernel for full parameters $\mathbf{K}_{\bm{\theta}}$, usually within the standard error of 5 runs using different random seeds. The performance of kernel logistic regression using $\mathbf{K}_{\bm{\theta}}$ is in turn close to full finetuning except for a few tasks including TREC, MNLI, SNLI, QNLI and MPQA, which are related to the prompt templates used. Next we directly compare the kernel matrices $\mathbf{K}_{\bm{\theta}}$, $\mathbf{K}_{\bm{\theta}_{R}}$ and $\mathbf{K}_{\bm{\theta}_{C}}$ for these few-shot learning problems directly. Figure 4 shows the empirical Neural Tangent Kernel values for the the task SST-2. More figures for the other tasks are available in Appendix E. This task is a two-class problem and hence their kernel matrices have 2x2 block structure. The values of the kernel entries are capped at 95-percentile for better visualization under heatmap. We can see that except for the magnitude of the entries in the kernel matrices, the patterns in the kernel matrices for the full parameter set $\mathbf{K}_{\bm{\theta}}$, 1-row set $\mathbf{K}_{\bm{\theta}_{R}}$ and 1-column set $\mathbf{K}_{\bm{\theta}_{C}}$ are extremely similar. More quantitatively, Table 6 shows the relative difference between the 1-row kernel $\mathbf{K}_{\bm{\theta}_{R}}$ and 1-column kernel $\mathbf{K}_{\bm{\theta}_{C}}$ with the full parameter kernel $\mathbf{K}_{\theta}$ after normalization in $\ell_{1}$ and $\ell_{2}$ norms by flattening the kernel matrices. For example, the relative difference for $\mathbf{K}_{\theta_{R}}$ is computed as
k-shot (single) Method SST-2 SST-5 MR CR MPQA Subj TREC
         16 Full FT 89.0(1.5) 44.6(1.4) 83.2(2.4) 93.3(0.2) 83.3(1.3) 88.5(2.6) 80.3(7.2)
$\mathbf{K}_{\bm{\theta}}$ 88.3(0.3) 43.6(2.2) 84.7(1.5) 93.2(0.9) 76.4(2.7) 88.6(1.3) 56.0(9.2)
$\mathbf{K}_{\bm{\theta}_{R}}$ 88.5(0.4) 42.9(1.9) 83.9(1.2) 93.2(0.5) 77.3(2.1) 85.8(1.2) 51.6(3.9)
$\mathbf{K}_{\bm{\theta}_{C}}$ 88.6(2.4) 42.4(1.9) 84.6(1.0) 93.2(0.5) 77.6(2.0) 85.9(1.2) 51.2(6.7)
         64 Full FT 89.7(0.4) 45.8(2.1) 85.6(1.1) 94.3(0.5) 84.8(0.8) 92.9(0.5) 93.2(1.0)
$\mathbf{K}_{\bm{\theta}}$ 89.2(1.0) 46.0(1.3) 86.4(0.6) 93.7(0.4) 81.2(0.9) 91.4(0.7) 77.8(2.3)
$\mathbf{K}_{\bm{\theta}_{R}}$ 89.5(0.5) 46.0(1.5) 86.4(0.6) 93.9(0.6) 81.6(0.7) 90.4(0.4) 70.4(1.6)
$\mathbf{K}_{\bm{\theta}_{C}}$ 89.5(0.6) 45.9(1.5) 86.4(0.4) 93.9(0.6) 81.5(0.5) 90.5(0.6) 70.7(2.5)
k-shot (pair) Method MNLI SNLI QNLI RTE MRPC QQP
         16 Full FT 59.2(2.7) 65.7(2.7) 62.1(3.1) 60.0(5.5) 73.9(2.7) 62.1(2.3)
         64 Full FT 68.7(1.7) 77.3(0.9) 72.8(2.2) 68.9(2.5) 82.8(1.2) 69.2(1.3)
(a) Full Parameter $\mathbf{K}_{\bm{\theta}}$
(b) 1-Row $\mathbf{K}_{\bm{\theta}_{R}}$
(c) 1-Column $\mathbf{K}_{\bm{\theta}_{C}}$
Figure 4: Neural Tangent Kernels for SST-2 on 16-shot training data.
16-shot (single) SST-2 SST-5 MR CR MPQA Subj TREC
16-shot (pair) MNLI SNLI QNLI RTE MRPC QQP
Table 6: Relative difference in kernels (compared to full parameter $\mathbf{K}_{\theta}$) on single-sentence and sentence-pair tasks.
6 Ablation Studies
Robustness of Row-Column Selection: In this study, we demonstrate the robustness of our row and column selection method through a detailed comparison of four selection strategies: Max, Min, Mixed, and Random. These strategies are applied to both rows and columns of the weight matrices. For the Min, Max, and Mixed selection strategies, we employ a scoring criterion used in the Wanda method (Sun et al., 2023), a simple yet effective pruning technique that requires only the forward pass. Pruning a neural network involves scoring the weights by importance (e.g., by the absolute values of weights), and then remove the least important ones. We can adopt these strategies to rank rows and columns by importance and evaluate the effect of finetuning on them. Given a weight matrix $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ and input feature activations $\mathbf{X}\in\mathbb{R}^{s\times d_{in}}$ from a length $s$ sequence, Wanda calculates the importance score $\mathbf{S}_{ij}$ of the weight $\mathbf{W}_{ij}$ as
$\mathbf{S}_{ij}=|\mathbf{W}_{ij}|\cdot\|\mathbf{X}_{\cdot j}\|_{2},$ (1)
@@ 6 Ablation Studies
Optimal Rank $r$ for RoCoFT : We investigate the impact of varying the rank $r$ on the performance of RoCoFT (Row and Column Fine-Tuning) and compare it with the widely used LoRA method within the RoBERTa-base attention block. We assess key metrics such as training time, accuracy, number of parameters, and memory consumption for each rank $r\in\{1,2,4,8,64\}$ using the SST2 dataset. The results are summarized in Table 7.
@@
7 Conclusions
Appendix A $\Delta\mathbf{W}$ Representation
@@
Appendix B Hyper-parameters for RoCoFT
The hyperparameters used in RoCoFT are provided in Table B.
@@ Appendix B Hyper-parameters for RoCoFT
Appendix C Environmental Setup and Implementation Details
@@
Appendix D RoCoFT with random weight selection
To test our hypothesis that finetuning LLMs can work as long as there are sufficient number of free parameters spread throughout the LLM model for training, we implement a version RoCoFT where instead of rows and columns, we randomly sample entries with probability $p$ from the weight matrices for updates and freeze the rest. Note that this method is not computationally efficient compared to updating only rows and columns and is only meant for ablation studies. From Tables 10 and 11 from we can see that updating random entries in the weight matrix is competitive with all other PEFT methods (we use $p=0.1$ and $p=0.01$ in these experiments). This gives further evidence that most good features are already acquired during pretraining and little learning is required during the finetuning stage.
LM # TTPs CoLA SST2 MRPC STS-B QQP MNLI QNLI RTE Avg.
RobertaBase 12.4M 63.15 94.96 88.03/89.18 90.57/90.07 89.29/86.94 87.22 92.60 80.01 85.72/89.62
RobertaLarge 35.5M 65.32 96.59 90.93/92.03 92.10/92.05 90.97/86.78 90.89 95.06 87.91 88.72/90.29
Table 10: RoBERTa models performance on GLUE tasks using 10% Random selection of trainable weights.
@@ Appendix D RoCoFT with random weight selection
Table 11: Accuracy comparison of commonsense and mathematical reasoning performance across different datasets using LLMs, with 1% of weights randomly selected for training.
Appendix E Additional Neural Tangent Kernel Results
Here we include additional results on our Neural Tangent Kernel experiments. Figure 6 shows the eigenvalue distribution of the full kernel $\mathbf{K}_{\bm{\theta}}$, 1-row kernel $\mathbf{K}_{\bm{\theta}_{R}}$ and 1-column kernel $\mathbf{K}_{\bm{\theta}_{C}}$ on different datasets. The eigenvalues are rescaled per dataset and we can see the eigenvalue distributions are very similar for the three NTK kernels. Table 12 shows the $\ell_{1}$ and $\ell_{2}$ norm difference between the kernel matrices of the 64-shot tasks, and the results are largely similar to the 16-shot results. The difference is mostly within 5-15%, but with smaller standard deviation than the 16-shot results over 5 random seeds. In Figure 7, we include a few more visualizations of the kernel matrices for the 16-shot tasks. We can see the three type of NTK matrices show very similar patterns across all tasks.
Figure 6: Eigenvalue spectrum of $\mathbf{K}_{{}_{\bm{\theta}}}$, $\mathbf{K}_{\bm{\theta}_{{}_{R}}}$, and $\mathbf{K}_{{}_{\bm{\theta}_{{}_{C}}}}$. The eigenvalues with respect to each task are scaled with $\beta_{{}_{task}}$ for representation.

### v3

@@ Appendix A Additional Experiments
Table 7: Results of DeBERTaV3-base on SQuAD v1.1, v2.0 benchmarks, reported using EM/F1 scores and BART-large on XSum and CNN/Daily Mail, reported using ROUGE metrics as ROUGE-1/ROUGE-2/ROUGE-L.
Appendix B More Ablation Study
@@
B.1 Optimal Rank $r$ for RoCoFT
We investigate the impact of varying the rank $r$ on the performance of RoCoFT (Row and Column Fine-Tuning) and compare it with the widely used LoRA method within the RoBERTa-base attention block. We assess key metrics such as training time, accuracy, number of parameters, and memory consumption for each rank $r\in\{1,2,4,8,64\}$ using the SST2 dataset. The results are summarized in Table8.
@@ B.1 Optimal Rank $r$ for RoCoFT
B.2 RoCoFT with random weight selection
To test our hypothesis that finetuning LLMs can work as long as there are sufficient number of free parameters spread throughout the LLM model for training, we implement a version RoCoFT where instead of rows and columns, we uniformly sample entries with probability $p$ from the weight matrices for updates and freeze the rest. Note that this method is not computationally efficient compared to updating only rows and columns and is only meant for ablation studies. From Tables LABEL:tab:roberta_glue_random and LABEL:table:LLM_results_random we can see that updating random entries in the weight matrix is competitive with all other PEFT methods (we use $p=0.1$ and $p=0.01$ in these experiments). This gives further evidence that most good features are already acquired during pretraining and little learning is required during the finetuning stage.
Appendix C Implementation
In Algorithm 1, we present a simplified PyTorch implementation of RoCoFT (row version). The main idea is to replace the linear layers in transformer model with the shown module, where $r$ rows are selected to be trainable weights and the remaining ones are frozen and converted to buffers (so that their gradients are not computed during backward pass). The forward function is simply the same as the original linear layer with the weights a concatenation of trainable and frozen weights. The version for columns are implemented similarly.
class RoCoFTRow(nn.Module):
    # inputs: F is the Linear Layer to be converted
    #         r is the rank (number of rows/columns to be selected)
    #         use_bias decides whether to train bias term
    def __init__(self, F, r, use_bias):
        # Set rows 1 to rank r as trainable weights
        self.trainable_W = nn.Parameter(F.weight[:r, :].clone())
        # Set rows r and above as non-trainable and move to buffer
        self.register_buffer(’non_trainable_W’, F.weight[r:, :].clone().detach())
        # Handle bias
        if F.bias is not None:
            self.bias = nn.Parameter(F.bias.clone().detach(), requires_grad=use_bias)
        else:
            self.bias = None
    def forward(self, x):
        full_weight = torch.cat([self.trainable_W, self.non_trainable_W], dim=1)
        out = torch.nn.functional.linear(x, full_weight, self.bias)
        return out
Algorithm 1 PyTorch pseudocode for replacing a linear layer in transformer model (row version)
Appendix D $\Delta\mathbf{W}$ Representation
@@
Appendix E Hyper-parameters for RoCoFT
@@
The hyperparameters used in RoCoFT are provided in Table 10.
Appendix F Environmental Setup and Implementation Details
@@
LM # TTPs CoLA SST2 MRPC STS-B QQP MNLI QNLI RTE
RobertaBase 12.4M 63.15 94.96 88.03/89.18 90.57/90.07 89.29/86.94 87.22 92.60 80.01
RobertaLarge 35.5M 65.32 96.59 90.93/92.03 92.10/92.05 90.97/86.78 90.89 95.06 87.91
Table 11: RoBERTa models performance on GLUE tasks using 10% random sampling of trainable parameters from each weight matrix ($p=0.1$).tab:roberta_glue_random
@@
Table 12: Accuracy comparison of commonsense and mathematical reasoning performance across different datasets using LLMs, using 1% random sampling of total trainable model parameters from each weight matrix ($p=0.01$).table:LLM_results_random
Appendix G Additional Neural Tangent Kernel Results
Here we include additional results on our Neural Tangent Kernel experiments. Figure LABEL:fig:spectrum shows the eigenvalue distribution of the full kernel $\mathbf{K}_{\bm{\theta}}$, 1-row kernel $\mathbf{K}_{\bm{\theta}_{R}}$ and 1-column kernel $\mathbf{K}_{\bm{\theta}_{C}}$ on different datasets. The eigenvalues are rescaled per dataset and we can see the eigenvalue distributions are very similar for the three NTK kernels. Table LABEL:tab:kernel_diff_single_k64 shows the $\ell_{1}$ and $\ell_{2}$ norm difference between the kernel matrices of the 64-shot tasks, and the results are largely similar to the 16-shot results. The difference is mostly within 5-15%, but with smaller standard deviation than the 16-shot results over 5 random seeds. In Figure G, we include a few more visualizations of the kernel matrices for the 16-shot tasks. We can see the three type of NTK matrices show very similar patterns across all tasks.
Figure 4: Eigenvalue spectrum of $\mathbf{K}_{{}_{\bm{\theta}}}$, $\mathbf{K}_{\bm{\theta}_{{}_{R}}}$, and $\mathbf{K}_{{}_{\bm{\theta}_{{}_{C}}}}$. The eigenvalues with respect to each task are scaled with $\beta_{{}_{task}}$ for better representation.fig:spectrum